Alyssa dropped the golf ball from the 5th floor of the hotel and hits the ground for 5 seconds later. How tall is a 5th floor hotel?


1
Expert's answer
2022-04-07T13:09:11-0400

The height of a 5th floor level

"h=gt^2/2=9.8*5^2/2=122.5\:\rm m"